OpenStudy (anonymous):

Suppose f(x)=4x-2 and g(x)=-2x+1. Find the value of f(0)/g(4)

OpenStudy (gorv):

f(0)=4*0-2=-2 g(4)=-2*4+1=-8+1=-7

OpenStudy (anonymous):

2\7

OpenStudy (gorv):

f(0)/g(4)=-2/-7=2/7

OpenStudy (anonymous):

thanks guys!